OpenStudy (anonymous):

the computer club went on a feild trip to a computer museum. The trip cost $240 to be paid equally by each club member. The day before the trip, 4 students decided not to go. This increased the cost by $2 dollars per student . How many students went to the computer museum?

OpenStudy (anonymous):

if you let x = initial number of students who went to the museum y = initial cost of trip per student 240 = xy (1) 240 = (x-4)(y+2) (2) solving to equation will lead solving x and y.... but we need x-4, the final number of students who went in museum

OpenStudy (anonymous):

@Orion1213 how do you solve the 2nd equation?

OpenStudy (anonymous):

you have 2 equations 1 and 2, and 2 unknowns... solving them simultaneously will lead you either to solve x or y first... then the other can be solve using either 1 or 2, once you plug-in the first one you have obtained...

OpenStudy (anonymous):

for equation 1... you can re-write it as...\[y=240/x\]and substitute or plug-in at equation 2 as...\[240=(x-4)(240/x +2)\]here you can solve the original number of students...

OpenStudy (anonymous):

hope you can make it from here...

OpenStudy (anonymous):

yeah. thank you